# Eigenvalue Gaps of the Laplacian of Random Graphs

### v1

@@ 5.1. Decomposition of vectors
  • Step 1: as $\mathbf{LCD}({\mathbf{v}}_{I})$ is small, ${\mathbf{v}}_{I}$ has structure. However, as we don’t have any information on ${\mathbf{v}}_{I^{c}}$, we will use the guide discussion in Subsection 2.5 to solve for ${\mathbf{v}}_{I^{c}}$ by conditioning on some parts of the matrix $\overline{L}_{n}$. We will do so by relying on Theorem 2.6, see (41).
  • Step 2: starting from (41), which roughly can be expressed as an event $\|M_{11}{\mathbf{v}}_{I}-{\mathbf{f}}\|_{2}$ being small for some fixed ${\mathbf{f}}$, we will work with the randomness of $M_{11}$, a principle minor of size $n_{0}\times n_{0}$. As this matrix is symmetric, we have to decompose into smaller rectangular blocks to see independent random rows and entries, see Figure 7. One major problem here is that the entries on the diagonal linearly depend on other entries of the rectangular blocks, and hence we have to use the $\rho_{L}$ notion from Theorem 5.1 rather than its $\rho$ counterpart. Because there are many shifts to deal with, we will start from the shift giving worst small ball probability (see (35)). This approach is significantly different from the regularized $\mathbf{LCD}$ structures used in [Ver14] to deal with random symmetric matrices.
Toward the second step, we will need to work with some more special form of ${\mathbf{v}}_{I}$ rather than knowing that its $\mathbf{LCD}$ is large. We first fix a subset $I$ of size $n_{0}$ in $[n]$. Divide $I$ into $k=\lfloor\sqrt{\log\log n}\rfloor$ intervals $I_{1},\dots,I_{k}$ (say of indices arranged consecutively in $I$), each of size approximately
@@
Given $A$, with sufficiently large $C$ (which is also used in the approximations of Subsection 2.4), and let
@@
Define  [^6] We note that $D({\mathbf{v}}_{I})$ does depend on the partition sets $I_{1},\dots,I_{k}$ and on $\kappa,\gamma^{\prime}$, but we bypass this dependency in the notation for simplicity.
$D({\mathbf{v}}_{I}):=\max_{j}\min_{a_{j}\in{\mathcal{A}}}\mathbf{LCD}_{\kappa,\gamma^{\prime}}(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}})=\max_{j}D_{j}.$ (35)
@@
Assume that $\min_{a\in{\mathcal{A}}}\mathbf{LCD}_{\kappa,\gamma^{\prime}}(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}})=D_{j}$, where $D_{j}=n^{{O(1)}}$. Then for any $a\in{\mathcal{A}}$, for any $r\geq 1/D_{j}$,
@@
By definition, for any $a\in{\mathcal{A}}$ we have $\mathbf{LCD}_{\kappa,\gamma^{\prime}}(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}})\geq D_{j}$. Hence by Theorem 3.7 (noting that under ${\mathcal{E}}_{non-gap}$ (36) holds), for any $r\geq 1/D_{j}\geq 1/\mathbf{LCD}_{\kappa,\gamma^{\prime}}(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}})$,
@@
Note that within this event, by definition we have every $D_{j}\ll D_{0}$, and hence we can $O(\kappa/D)$-approximate all the subvectors ${\mathbf{v}}_{j}-a_{j}\mathbf{1}$ (for some $a_{j}\in{\mathcal{A}}$). Note that the norm of each of the vector ${\mathbf{v}}_{I_{j}}-a_{j}\mathbf{1}$ is at least $c_{n}$ by (36), so by Lemma 3.8, by gluing the nets of the subvectors together, we obtain an $O(\kappa k/D_{0})$-net ${\mathcal{N}}_{I}$ of size $(C_{0}c_{n}^{-1}D_{0}/\sqrt{m})^{km}|{\mathcal{A}}|^{k}$ for the vectors ${\mathbf{v}}_{I}$, where (37) holds.
Now for a fixed ${\mathbf{v}}_{I}$ in ${\mathcal{N}}_{I}$, consider the event that there exists ${\mathbf{v}}_{I^{c}}\in{\mathbb{R}}^{n-n_{0}}$ (of norm of order 1, more specifically ${\mathbf{v}}_{I^{c}}\in\frac{1}{n^{C+1}}{\mathbb{Z}}^{n-n_{0}}$) so that
@@
In particularly, $\|M_{21}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}\|_{2}\leq\varepsilon$.
@@
$k_{0}:=d_{n}(n-n_{0}),\mbox{ where, say }d_{n}=(\frac{1}{\log n})^{2}.$ (40)
@@
Then from $\|M_{21}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}\|_{2}\leq\varepsilon$ of (39), after being applied by $M_{22}^{\prime}$, we deduce that
@@
In other words, if we fix a realization of $M_{21}$ (and hence $M_{12}$) and $M_{22}$, and if we fix $T$ and a choice of ${\mathbf{v}}_{I^{c}}|_{T^{c}}$, then ${\mathbf{v}}_{I^{c}}|_{T}$ lies within $M_{22}^{\prime}M_{21}{\mathbf{v}}_{I}+M_{22}^{\prime}M_{22}|_{T^{c}}{\mathbf{v}}_{I^{c}}|_{T^{c}}$ of distance at most $\varepsilon/d_{n}\sqrt{n}$. So if we write $M_{12}{\mathbf{v}}_{I^{c}}$ in the form $M_{12}^{\prime}{\mathbf{v}}_{I^{c}}|_{T^{c}}+M_{12}^{\prime\prime}{\mathbf{v}}_{I^{c}}|_{T}$ then we have
@@
where we recall that under ${\mathcal{E}}_{iid-norm}$, $\|M_{12}\|_{2}=O(\sqrt{n})$. Together with the first half of (39), $\|M_{11}{\mathbf{v}}_{I}+M_{12}{\mathbf{v}}_{I^{c}}\|_{2}\leq\varepsilon$, by triangle inequality
@@
where $F$ is a deterministic matrix, and where for short we set
@@
for some deterministic vector ${\mathbf{w}}$, where the randomness now is on the entries of $(M_{11})_{(I\backslash I_{j_{0}})}$, which are all independent.
Next, for each $i\in I\backslash I_{j_{0}}$ where $|v_{i}|\leq n^{{-1/2}}\log^{2}n$ we consider the event $|{\mathbf{r}}_{i}(M_{11})_{(I\backslash I_{j_{0}})\times I_{j_{0}}}\cdot{\mathbf{v}}_{I_{j_{0}}}-w_{i}|=O(\rho)$, which has the form
$|x_{1}(v_{i_{1}}-v_{i})+\dots+x_{m}(v_{i_{m}}-v_{i})-f|=O(\rho)$
where $x_{1},\dots,x_{m}$ are independent Bernoulli of parameter $p$, and where we recall that $km=n_{0}$ (with $k=\lfloor\sqrt{\log\log n}\rfloor$ and $n/k\leq n_{0}\leq n$), and $i_{1},\dots,i_{m}$ are the elements of $I_{j_{0}}$ and $f$ might depend on $v_{i_{j}}$ but not on the $x_{i}$.
We note that as $\sum_{i\in I_{j_{0}}}v_{i}^{2}\leq 1$, the set indices $i$ where $|v_{i}|\leq n^{{-1/2}}\log^{2}n$ is at least $n_{0}-m-n/\log^{4}n$. By Claim 6.2, as $D_{j_{0}}=D({\mathbf{v}}_{I})\in[D_{0},2D_{0}]$ and as $\rho\geq 1/D_{0}$,
@@
By Lemma 3.3, we thus have
@@
Summing over $(C_{0}c_{n}^{-1}D_{0}/\sqrt{m})^{n_{0}}|{\mathcal{A}}|^{k}$ choices for the vector ${\mathbf{v}}_{I}$, over $(n^{C+1})^{d_{n}(n-n_{0})}\times\binom{n-n_{0}}{d_{n}(n-n_{0})}$ choices for the vector ${\mathbf{v}}_{I^{c}}|_{T^{c}}$, and over $\binom{n}{n_{0}}$ choices for $I$ we obtain the final union bound
$(\rho\log^{{O(1)}}n)^{n_{0}-m-n/\log^{4}n}\times(C_{0}c_{n}^{-1}D_{0}/\sqrt{m})^{n_{0}}|{\mathcal{A}}|^{k}\times(n^{C+1})^{d_{n}(n-n_{0})}\times\binom{n-n_{0}}{d_{n}(n-n_{0})}\times\binom{n-n_{0}}{d_{n}(n-n_{0})}\times\binom{n}{n_{0}}$
@@
Proof.
(of Therem 2.1) Let $\overline{L}_{n}^{\prime}$ be obtained from $\overline{L}_{n}$ by changing the last two columns (and rows respectively) ${\mathbf{c}}_{n-1},{\mathbf{c}}_{n})$ of $\overline{L}_{n}$ to $({\mathbf{c}}_{n-1}+{\mathbf{c}}_{n})/\sqrt{2}$ and $(-{\mathbf{c}}_{n-1}+{\mathbf{c}}_{n})/\sqrt{2}$. We observe the followings.
@@
Our next lemma is that
@@
The conclusion of Theorem 2.3 and 2.4 holds for the egivenvectors of all principles $\overline{L}^{\prime}_{n-1}$ of $\overline{L}^{\prime}_{n}$ of size $(n-1)\times(n-1)$.
@@
The proof of this result is identical to the proofs of Theorems 2.3 and 2.4; therefore we omit the details. ∎
Let ${\mathbf{u}}=\left(\begin{array}[]{c}{\mathbf{w}}\\ b\end{array}\right)$ be an eigenvector of $\overline{L}^{\prime}_{n}$, where ${\mathbf{w}}\in{\mathbb{R}}^{n-1}$, with eigenvalue $\lambda_{i}(\overline{L}^{\prime}_{n})$. By Theorem 2.3 and Claim 3.4 we might assume that $|b|\geq n^{{-1/2-o(1)}}$ (as these results imply that with probability $1-n^{-\omega(1)}$ the set of indices where $|b|\geq n^{{-1/2-o(1)}}$ is of size $\Theta(n)$).
@@
$\overline{L}^{\prime}_{n}=\left(\begin{array}[]{cc}\overline{L}^{\prime}_{n-1}&{\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})\\ {\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})^{T}&f\end{array}\right).$
where ${\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})$ is a $(n-1)\times 1$ vector  [^7] We slightly abused our standard notation, here ${\mathbf{c}}_{n}$ is not exactly the last column of $\overline{L}^{\prime}_{n}$ but only a subvector. and $f\in{\mathbb{R}}$. Argue as in the discussion following Theorem 2.1, let ${\mathbf{v}}$ be a unit vector of $\overline{L}^{\prime}_{n-1}$ corresponding to $\lambda_{i}(\overline{L}^{\prime}_{n-1})$, then we have
@@
We next conclude with a proof of the non-centered model, where we recall that in our notation, the real eigenvalues of a symmetrix matrix $M_{n}$ are arranged as $\lambda_{1}(M_{n})\leq\dots\leq\lambda_{n}(M_{n})$.
Proof.
(of Therem 1.2) Now that we have established the result for $\overline{L}_{n}$, we translate the result back to $L_{n}$. Assume that the spectral decomposition of $\overline{L}_{n}$ is (with $\lambda_{1}\leq\dots\leq\lambda_{n}$)
@@
where one of the eigenvalue is $\lambda_{i_{0}}=0$, with ${\mathbf{v}}_{i_{0}}=\overline{\mathbf{1}}=\mathbf{1}/\sqrt{n}$, and where the eigenvectors ${\mathbf{v}}_{i}$ forms an orthonormal basis. Note that we have given effective gaps between the $\lambda_{i}$.
Observe that ${\mathbf{E}}L_{n}=-pJ_{n}+pnI_{n}$ where $J_{n}$ is the $n\times n$ matrix of ones. So
@@
Since $L_{n}$ is positive semidefinite, the zero eigenvalue of $L_{n}$ is the smallest eigenvalue $\lambda_{1}(L_{n})$, with eigenvector $\overline{\mathbf{1}}$  [^8] Also, $L_{n}$ and $\overline{L}_{n}$ have the same set of eigenvectors.. So for Therem 1.2, it suffices to bound the gap of the two smallest eigenvalues of $L_{n}$, $\lambda_{2}(L_{n})-\lambda_{1}(L_{n})=\lambda_{2}(L_{n})$.
@@
(of Therem 2.7) As $\overline{L}_{n}$ and $L_{n}$ have the same set of eigenvectors, it suffices to work with $\overline{L}_{n}$. We consider the decomposition
$\overline{L}_{n}=\left(\begin{array}[]{ccc}\overline{L}_{n-2}&{\mathbf{c}}_{n-1}(\overline{L}_{n})&{\mathbf{c}}_{n}(\overline{L}_{n})\\ {\mathbf{c}}_{n-1}(\overline{L}_{n})^{T}&f&g\\ {\mathbf{c}}_{n}(\overline{L}_{n})^{T}&g&h\end{array}\right).$
@@
In other words, ${\mathbf{w}}$ is an approximated vector of $\overline{L}_{n-2}$.
@@
$|{\mathbf{c}}_{n}(\overline{L}_{n})^{T}{\mathbf{w}}|\leq|ga|+|hb|+|\lambda_{i}(\overline{L}_{n})||b|$
@@
The rest of the proof is almost identical to the proof of Therem 2.1. Let $I_{n-2}$ be the set of vertices $v_{1},\dots,v_{n-2}$ which has exactly one neighbor with $v_{n}$ and $v_{n-1}$. Then the neighbor switching process changes the coordinates over $I_{n-2}$ of ${\mathbf{c}}_{n}(\overline{L}_{n})$ randomly and independently. Furthermore, with probability $1-\exp(-\Theta(n))$, we know that
@@
We can then take a union bound over the $\binom{n}{2}$ possible sets of pairs of coordinates.

### v2

@@ 5.1. Decomposition of vectors
  • Step 1: If $\mathbf{LCD}({\mathbf{v}}_{I})$ is small, ${\mathbf{v}}_{I}$ has structure. However, as we don’t have any information on ${\mathbf{v}}_{I^{c}}$, we will use as a guide the discussion in Subsection 2.5 to solve for ${\mathbf{v}}_{I^{c}}$ by conditioning on some parts of the matrix $\overline{L}_{n}$. We will do so by relying on Theorem 2.6, see (43).
  • Step 2: Starting from (43), which roughly can be expressed as an event $\|M_{11}{\mathbf{v}}_{I}-{\mathbf{f}}\|_{2}$ being small for some fixed ${\mathbf{f}}$, we will work with the randomness of $M_{11}$, a principal minor of size $n_{0}\times n_{0}$. As this matrix is symmetric, we have to decompose it into smaller rectangular blocks to see independent random rows and entries, see Figure 7. One major problem here is that the entries on the diagonal depend on other entries of the rectangular blocks, and hence we have to use the $\rho_{L}$ notion from Theorem 5.1 rather than its $\rho$ counterpart. Because there are many shifts to deal with, we will start from the shift giving the worst small ball probability (see (37)). This approach is significantly different from the regularized $\mathbf{LCD}$ structures used in [66] to deal with random symmetric matrices.
Toward the second step, we will need to work with some more special form of ${\mathbf{v}}_{I}$, rather than knowing that its $\mathbf{LCD}$ is large. We first fix a subset $I$ of size $n_{0}$ in $[n]$. Divide $I$ into $k=\lfloor\sqrt{\log\log n}\rfloor$ intervals $I_{1},\dots,I_{k}$ (for instance, of indices arranged consecutively in $I$), each of size approximately
@@
Given $A$ and sufficiently large $C$ (which is also used in the approximations of Subsection 2.4), let
@@
Define  [^7] We note that $D({\mathbf{v}}_{I})$ does depend on the partition sets $I_{1},\dots,I_{k}$ and on $\kappa,\gamma^{\prime}$, but we suppress this dependency in the notation for simplicity.
$D({\mathbf{v}}_{I}):=\max_{j}\min_{a_{j}\in{\mathcal{A}}}\mathbf{LCD}_{\kappa,\gamma^{\prime}}\left(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}}\right)=\max_{j}D_{j}.$ (37)
@@
Assume that $\min_{a\in{\mathcal{A}}}\mathbf{LCD}_{\kappa,\gamma^{\prime}}\left(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}}\right)=D_{j}$, where $D_{j}=n^{{O(1)}}$. Then for any $a\in{\mathcal{A}}$ and $r\geq 1/D_{j}$,
@@
By definition, for any $a\in{\mathcal{A}}$, we have $\mathbf{LCD}_{\kappa,\gamma^{\prime}}\left(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}}\right)\geq D_{j}$. Hence by Theorem 3.7 (noting that under ${\mathcal{E}}_{non-gap}$, (38) holds), for any $r\geq 1/D_{j}\geq 1/\mathbf{LCD}_{\kappa,\gamma^{\prime}}\left(\frac{{\mathbf{v}}_{I_{j}}-a\mathbf{1}}{\|{\mathbf{v}}_{I_{j}}-a\mathbf{1}\|_{2}}\right)$,
@@
Note that within this event, by definition we have every $D_{j}\ll D_{0}$, and hence we can $O(\kappa/D)$-approximate all the subvectors ${\mathbf{v}}_{j}-a_{j}\mathbf{1}$ (for some $a_{j}\in{\mathcal{A}}$). Note that the norm of each of the vector ${\mathbf{v}}_{I_{j}}-a_{j}\mathbf{1}$ is at least $c_{n}$ by (38), so by Lemma 3.8, by gluing the nets of the subvectors together, we obtain a $O(\kappa k/D_{0})$-net ${\mathcal{N}}_{I}$ of size $(C_{0}c_{n}^{-1}D_{0}/\sqrt{m})^{km}|{\mathcal{A}}|^{k}$ for the vectors ${\mathbf{v}}_{I}$ where (39) holds.
Now, for a fixed ${\mathbf{v}}_{I}$ in ${\mathcal{N}}_{I}$, consider the event that there exists ${\mathbf{v}}_{I^{c}}\in{\mathbb{R}}^{n-n_{0}}$ (of norm of order 1, more specifically ${\mathbf{v}}_{I^{c}}\in\frac{1}{n^{C+1}}{\mathbb{Z}}^{n-n_{0}}$) so that
@@
In particular, $\|M_{21}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}\|_{2}\leq\varepsilon$.
@@
$k_{0}:=d_{n}(n-n_{0}),\mbox{ where }d_{n}=\left(\frac{1}{\log n}\right)^{2}.$ (42)
@@
Then, from the equation $\|M_{21}{\mathbf{v}}_{I}+M_{22}{\mathbf{v}}_{I^{c}}\|_{2}\leq\varepsilon$ coming from (41), we apply $M_{22}^{\prime}$, which yields
@@
In other words, if we fix a realization of $M_{21}$ (and hence $M_{12}$) and $M_{22}$, and if we fix $T$ and a choice of ${\mathbf{v}}_{I^{c}}|_{T^{c}}$, then ${\mathbf{v}}_{I^{c}}|_{T}$ lies within a distance of at most $\varepsilon/d_{n}\sqrt{n}$ to $-M_{22}^{\prime}M_{21}{\mathbf{v}}_{I}-M_{22}^{\prime}M_{22}|_{T^{c}}{\mathbf{v}}_{I^{c}}|_{T^{c}}$. Using this approximation in our other equation yields
@@
where we recall that under ${\mathcal{E}}_{i.i.d.-norm}$, $\|M_{12}\|_{2}=O(\sqrt{n})$. Together with the first half of (41), $\|M_{11}{\mathbf{v}}_{I}+M_{12}{\mathbf{v}}_{I^{c}}\|_{2}\leq\varepsilon$, by the triangle inequality we have
@@
where $F$ is a deterministic matrix, and where for brevity we write
@@
for some deterministic vector ${\mathbf{w}}$, where the randomness now is from the entries of $(M_{11})_{(I\backslash I_{j_{0}})}$, which are all independent.
Next, for each $i\in I\backslash I_{j_{0}}$ where $|v_{i}|\leq n^{{-1/2}}\log^{2}n$, we consider the event $|{\mathbf{r}}_{i}(M_{11})_{(I\backslash I_{j_{0}})\times I_{j_{0}}}\cdot{\mathbf{v}}_{I_{j_{0}}}-w_{i}|=O(\rho)$, which has the form
$|x_{1}(v_{i_{1}}-v_{i})+\dots+x_{m}(v_{i_{m}}-v_{i})-f|=O(\rho),$
where $x_{1},\dots,x_{m}$ are independent Bernoulli of parameter $p$, and where we recall that $km=n_{0}$ (with $k=\lfloor\sqrt{\log\log n}\rfloor$ and $n/k\leq n_{0}\leq n$), $i_{1},\dots,i_{m}$ are the elements of $I_{j_{0}}$, and $f$ might depend on $v_{i_{j}}$ but not on the $x_{i}$.
We note that as $\sum_{i\in I_{j_{0}}}v_{i}^{2}\leq 1$, the set indices $i$ where $|v_{i}|\leq n^{{-1/2}}\log^{2}n$ has cardinality at least $n_{0}-m-n/\log^{4}n$. By Claim 6.2, as $D_{j_{0}}=D({\mathbf{v}}_{I})\in[D_{0},2D_{0}]$ and as $\rho\geq 1/D_{0}$,
@@
Therefore, by Lemma 3.3, we have
@@
Summing over $(C_{0}c_{n}^{-1}D_{0}/\sqrt{m})^{n_{0}}|{\mathcal{A}}|^{k}$ choices for the vector ${\mathbf{v}}_{I}$, over $(n^{C+1})^{d_{n}(n-n_{0})}\times\binom{n-n_{0}}{d_{n}(n-n_{0})}$ choices for the vector ${\mathbf{v}}_{I^{c}}|_{T^{c}}$, and over $\binom{n}{n_{0}}$ choices for $I$, we obtain the final union bound
$(\rho\log^{{O(1)}}n)^{n_{0}-m-n/\log^{4}n}(C_{0}c_{n}^{-1}D_{0}/\sqrt{m})^{n_{0}}|{\mathcal{A}}|^{k}(n^{C+1})^{d_{n}(n-n_{0})}\binom{n-n_{0}}{d_{n}(n-n_{0})}\binom{n-n_{0}}{d_{n}(n-n_{0})}\binom{n}{n_{0}}$
@@
Proof of Theorem 2.1.
Let $\overline{L}_{n}^{\prime}$ be obtained from $\overline{L}_{n}$ by changing the last two columns (and rows respectively) ${\mathbf{c}}_{n-1},{\mathbf{c}}_{n})$ of $\overline{L}_{n}$ to $({\mathbf{c}}_{n-1}+{\mathbf{c}}_{n})/\sqrt{2}$ and $(-{\mathbf{c}}_{n-1}+{\mathbf{c}}_{n})/\sqrt{2}$. We observe the followings.
@@
Our next lemma is the following.
@@
The conclusion of Theorem 2.3 and 2.4 holds for the eigenvectors of all principals $\overline{L}^{\prime}_{n-1}$ of $\overline{L}^{\prime}_{n}$ of size $(n-1)\times(n-1)$.
@@
The proof of this result is identical to the proofs of Theorems 2.3 and 2.4. Therefore, we omit the details. ∎
Let ${\mathbf{u}}=\left(\begin{array}[]{c}{\mathbf{w}}\\ b\end{array}\right)$ be an eigenvector of $\overline{L}^{\prime}_{n}$ with associated eigenvalue $\lambda_{i}(\overline{L}^{\prime}_{n})$, where ${\mathbf{w}}\in{\mathbb{R}}^{n-1}$. By Theorem 2.3 and Claim 3.4, we assume that $|b|\geq n^{{-1/2-o(1)}}$ (as these results imply that with probability $1-n^{-\omega(1)}$ the set of indices where $|b|\geq n^{{-1/2-o(1)}}$ is of size $\Theta(n)$).
@@
$\overline{L}^{\prime}_{n}=\left(\begin{array}[]{cc}\overline{L}^{\prime}_{n-1}&{\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})\\ {\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})^{T}&f\end{array}\right),$
where ${\mathbf{c}}_{n}(\overline{L}^{\prime}_{n})$ is a $(n-1)\times 1$ vector  [^8] We slightly abused our standard notation, here ${\mathbf{c}}_{n}$ is not exactly the last column of $\overline{L}^{\prime}_{n}$ but only a subvector. and $f\in{\mathbb{R}}$. Arguing as in the discussion following Theorem 2.1, let ${\mathbf{v}}$ be a unit vector of $\overline{L}^{\prime}_{n-1}$ corresponding to $\lambda_{i}(\overline{L}^{\prime}_{n-1})$. We have
@@
We next conclude with a proof of the non-centered model, where we recall that in our notation, the real eigenvalues of a symmetric matrix $M_{n}$ are arranged as $\lambda_{1}(M_{n})\leq\dots\leq\lambda_{n}(M_{n})$.
Proof of Theorem 1.2.
Now that we have established the result for $\overline{L}_{n}$, we translate the result back to $L_{n}$. Assume that the spectral decomposition of $\overline{L}_{n}$ is (with $\lambda_{1}\leq\dots\leq\lambda_{n}$)
@@
where one of the eigenvalues is $\lambda_{i_{0}}=0$, with ${\mathbf{v}}_{i_{0}}=\overline{\mathbf{1}}=\mathbf{1}/\sqrt{n}$, and where the eigenvectors ${\mathbf{v}}_{i}$ form an orthonormal basis. Note that we have given effective gaps between the $\lambda_{i}$.
Observe that ${\mathbf{E}}L_{n}=-pJ_{n}+pnI_{n}$, where $J_{n}$ is the $n\times n$ matrix of ones. So
@@
Since $L_{n}$ is positive semidefinite, the zero eigenvalue of $L_{n}$ is the smallest eigenvalue $\lambda_{1}(L_{n})$, with eigenvector $\overline{\mathbf{1}}$  [^9] Also, $L_{n}$ and $\overline{L}_{n}$ have the same set of eigenvectors.. So for Theorem 1.2, it suffices to bound the gap of the two smallest eigenvalues of $L_{n}$, $\lambda_{2}(L_{n})-\lambda_{1}(L_{n})=\lambda_{2}(L_{n})$.
@@
(of Theorem 2.7) As $\overline{L}_{n}$ and $L_{n}$ have the same set of eigenvectors, it suffices to work with $\overline{L}_{n}$. We consider the decomposition
$\overline{L}_{n}=\left(\begin{array}[]{ccc}\overline{L}_{n-2}&{\mathbf{c}}_{n-1}(\overline{L}_{n})&{\mathbf{c}}_{n}(\overline{L}_{n})\\ {\mathbf{c}}_{n-1}(\overline{L}_{n})^{T}&f&g\\ {\mathbf{c}}_{n}(\overline{L}_{n})^{T}&g&h\end{array}\right),$
@@
In other words, ${\mathbf{w}}$ is an approximate eigenvector of $\overline{L}_{n-2}$.
@@
$|{\mathbf{c}}_{n}(\overline{L}_{n})^{T}{\mathbf{w}}|\leq|ga|+|hb|+|\lambda_{i}(\overline{L}_{n})||b|,$
@@
The rest of the proof is almost identical to the proof of Theorem 2.1. Let $I_{n-2}$ be the set of vertices $v_{1},\dots,v_{n-2}$ which has exactly one neighbor with $v_{n}$ and $v_{n-1}$. Then the neighbor switching process changes the coordinates over $I_{n-2}$ of ${\mathbf{c}}_{n}(\overline{L}_{n})$ randomly and independently. Furthermore, with probability $1-\exp(-\Theta(n))$, we know that
@@
We can then take a union bound over the $\binom{n}{2}$ possible sets of pairs of coordinates. ∎